Quick Computing installed its previous generation of computer chip manufacturing equipment 3 years ago. Some of that older equipment will become unnecessary when the company goes into production of its new product. The obsolete equipment, which originally cost $41.50 million, has been depreciated straight-line over an assumed tax life of 5 years, but it can be sold now for $18.30 million. The firm’s tax rate is 30%. What is the after-tax cash flow from the sale of the equipment? (Enter your answer in millions rounded to 1 decimal place.)
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(41.5/5)=$8.3 million/year

Hence book value as on date of sale=Cost-Accumulated Depreciation

=41.5-(8.3*3)=16.6 million

Hence gain on sale=(18.3-16.6)=1.7 million

Hence after-tax cash flow=Sale proceeds-(Tax rate*Gain on sale)

=18.3-(1.7*0.3) million

=$17.8 million(Approx).